OpenStudy (anonymous):

Does anyone know how to solve this: y′′ − 2y′ − 3y = cos(2x)

OpenStudy (amistre64):

yes

OpenStudy (amistre64):

determine the homogenous part, then that can be used to determine the rest of it with either a wronskian approach or a parameters approach. They even start you off in this subject with a "table" approach

OpenStudy (anonymous):

Thank you. My lecturer did not really provide any good lecture material. So it's a bit hard to understand the mathematical terms that you used lol

OpenStudy (amistre64):

can you solve a quadratic equation?

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

some background then: assume e^(rx) is a solution y = e^(rx) y' = r e^(rx) y'' = r^2 e^(rx) sub it in for the homogenous part y′′ − 2y′ − 3y = 0 r^2 e^(rx) -2r e^(rx) -3e^(rx) = 0 factor out the e^(rx) e^(rx) (r^2 -2r -3) = 0 solve the quadratic part of it for r

OpenStudy (anonymous):

Hey thanks. I would be able to solve it if the cos(2x) wasn't there. Don't know what to do with the cos(2x).

OpenStudy (amistre64):

parameters is the long version, the wronskian is the shortcut. but its good to use parameters to understand why the wronskian is working

OpenStudy (amistre64):

im most likely confusing undetermined coeffs with parameters ... just too many words to keep track of lol

OpenStudy (anonymous):

Oh okay. I'm really sorry. I'm probably annoying you, but Ive never heard of anything called wronskian

OpenStudy (amistre64):

its a guys name, its related to the Cramer rule, another guys name. Its a matrix solution once you know the homogenous parts,

OpenStudy (anonymous):

Oh right. Hmm so do I have to find out how to solve this problem using wronskian?

OpenStudy (amistre64):

its folded away neatly into some randomized corner of me brain .... yes

OpenStudy (amistre64):

but lets solve the homogenous part; what do you get for "r" ?

OpenStudy (anonymous):

r=3 and r=-1

OpenStudy (amistre64):

then: y1 = A e^(3x), y2= B 2^(-x) y = yh + yp yh = y1+y2 = A e^(3x) + B e^(-x) ; is part of the solution yp is some function of x that we will now attempt to determine

OpenStudy (amistre64):

assume A and B are functions of x yp = A e^(3x) + B e^(-x) yp' = 3A e^(3x) - B e^(-x) + A' e^(3x) + B' e^(-x) , let A' and B' parts = 0 to account for the homogenous stuff yp'' = 9A e^(3x) + B e^(-x) + 3A' e^(3x) - B' e^(-x) now we can apply this into our setup

OpenStudy (amistre64):

y′′ − 2y′ − 3y = cos(2x) -3yp = -3A e^(3x) -3B e^(-x) -2yp' = -6A e^(3x) + 2B e^(-x) yp'' = 9A e^(3x) + B e^(-x) + 3A' e^(3x) - B' e^(-x) ------------------------------------------------- cos(2x) = 0 + 0 + 3A' e^(3x) - B' e^(-x)

OpenStudy (amistre64):

we now have 2 equations in 2 unknowns A' e^(3x) + B' e^(-x) = 0 A' 3e^(3x) - B' e^(-x) = cos(2x) solve for A' and B'

OpenStudy (anonymous):

Wow this is so long. I need to go through it again to make sure I understand it. Thank you so much!!! =D

OpenStudy (anonymous):

learn something new every day! or try to

OpenStudy (anonymous):

Yeah lol. Maths isn't easy at all!! I will do lots of research after my exams

OpenStudy (amistre64):

the Wronskian just assumes that you have already done a hundred of these and are at the end :)\[\begin{vmatrix} W_x&W&W_y\\ e^{3x}&0&e^{-x}\\ 3e^{3x}&cos(2x)&-e^{-x}\end{vmatrix}\] the parts needed to apply the Wronskian is the "determinant" of the matrix solution to this cramerized setup

OpenStudy (anonymous):

Oh that's quite interesting. I will look into that in more detail. Thanks once again!!

OpenStudy (amistre64):

youre welcome, if you havent come across this stuff yet, then its a good primer for whats in the next few chapters lol

OpenStudy (anonymous):

Lol thanks for the advice. Will prepare myself for next year then :)

OpenStudy (anonymous):

I did not get the right answer =(

OpenStudy (anonymous):

\[Ae ^{-x}+Be ^{3x}-\frac{ 1 }{ 65 }[7\cos(2x)+4\sin(2x)]\] this is what I should get "/

OpenStudy (amistre64):

A' e^(3x) + B' e^(-x) = 0 A' 3e^(3x) - B' e^(-x) = cos(2x) ---------------------------- A' 4e^(3x) = cos(2x) A' = cos(2x) e^(-3x) \[A=\int cos(2x) e^{-3x}=\frac1{13}e^{-3x}(2sin(2x)-3cos(2x))\] ................................................................. A' e^(3x) + B' e^(-x) = 0 -A' 3e^(3x) + B' e^(-x) = -cos(2x) ---------------------------- B' 2e^(-x) = -cos(2x) B' = -2cos(2x) e^(x) \[B=\int -2cos(2x) e^{x}=-\frac25 e^{x}(2sin(2x)+3cos(2x))\] \[y=ae^{3x}+be^{-x}+\frac1{13}e^{-3x}e^{3x}(2sin(2x)-3cos(2x))-\frac25 e^{x}e^{-x}(2sin(2x)+3cos(2x))\] \[y=ae^{3x}+be^{-x}+\frac1{13}(2sin(2x)-3cos(2x))-\frac25 (2sin(2x)+3cos(2x))\] \[y=ae^{3x}+be^{-x}+\frac2{13}sin(2x)-cos(2x)-\frac45sin(2x)-\frac65cos(2x)\] \[y=ae^{3x}+be^{-x}+\frac{42}{65}sin(2x)-\frac{11}5cos(2x)\] ... as long as im not making any simple arithmatic erros, it does look to be headed in that direction

OpenStudy (amistre64):

forgot to distribute the 1/13 to the -3cos(2x) ....

OpenStudy (anonymous):

Yeah that does look better. You know where you're trying to find A'. Where does the 4 go?

OpenStudy (amistre64):

... simple algebra mistakes is all, need to do a divide by 4, and a divide by 2 on B that might help align this stuff betterly

OpenStudy (amistre64):

A' = 1/4 e^(3x) cos(2x) B' = -1/2 e^(x) cos(2x) then integrate ....

OpenStudy (amistre64):

ive got to focus on my paying job at the moment, ill run thru what ive posted later to see what other mistakes might have crept in

OpenStudy (anonymous):

Aww That's so nice of you. Thank you once again.

OpenStudy (amistre64):

yeah, i made a few basic errors while trying to type all that in the latex code :) A' e^(3x) + B' e^(-x) = 0 A' 3e^(3x) - B' e^(-x) = cos(2x) A' = 1/4 e^(-3x) cos(2x) B' = -1/4 e^(x) cos(2x) A = 1/52 e^(-3x) (2 sin(2x)-3 cos(2x)) B = -1/20 e^x (2 sin(2x)+cos(2x)) 1/52 (2 sin(2x)-3 cos(2x)) -1/20 (2 sin(2x)+cos(2x)) which simplifies to: -1/65 (7cos(2x)+4sin(2x))

OpenStudy (anonymous):

Hey thanks! Wow you are a great helper!!

OpenStudy (amistre64):

depending on your level of integrating skill, this might be a bit difficult integrating A' and B' A more practical method is called "undetermined coefficients", and its pretty much a trial and error approach. we have already determined the yh parts that goes to zero, so we would need to determine a suitable yp since we want this to end up equaling cos(2x), make a guess at yp; yp = Acos(2x) + Bsin(2x), this time A and B are considered numbers, and not functions

OpenStudy (anonymous):

Hey yeah I know the second method

OpenStudy (anonymous):

Did some research and found out how to do it. However my mate told me that it doesn't always work. So I got quite stressed as an equation might come up where I cant use that method "/

OpenStudy (amistre64):

yp = A cos(2x) + B sin(2x) yp' = -2A sin(2x) + 2B cos(2x) yp'' = -4A cos(2x) - 4B sin(2x) plug them into our: y'' -2y' -3y = cos(2x) -3yp = -3A cos(2x) -3B sin(2x) -2yp' = 4A sin(2x) -4B cos(2x) yp'' = -4A cos(2x) - 4B sin(2x) ------------------------------ cos(2x) = (-7A-4B) cos(2x) + (-7B+4A) sin(2x) this only pans out if: 4A-7B = 0 and: -7A-4B = 1 this is again a system of equations in 2 unknowns, but no integration is needed, just algebra

OpenStudy (anonymous):

come up in the exam*

OpenStudy (amistre64):

true it doesnt always pan out, but it tends to be shorter and worth the effort to try :) there are tables printed of the most appropriate guesses for yp

OpenStudy (anonymous):

Ohh okay. I'm scared now. I do get ur method BUT I find it really hard to integrate A' and B'

OpenStudy (amistre64):

i found it a bit tedious to integrate it by hand and try to simplify it as well. Which is why i decided to suggest the undertermined coeff method:) most likely if you havent done the variation of parameters and wronskian, then you will only be asked diffyQs that relate to the best guess table

OpenStudy (anonymous):

Yeah I haven't done that. I hope you're right. Fingers crossed. I feel better now. Thank you ^_^

OpenStudy (amistre64):

good luck :)

OpenStudy (anonymous):

Thank you :D